# A case for specialisation in non-human entities

### v1

@@ 4 Why specialise?
4.1 Large models are more vulnerable
In the context of artificial intelligence, generality is obtained by training large models on massive web data crawls. Indeed, ever larger models and datasets have been empirically observed to yield ever more spectacular performances [BMR+20, KMH+20b, DJP+24]. However, these high performances seem inevitably entangled with a number of security issues [EMFG+22, OV23], like privacy violations, jailbreaking and data poisoning. Below we detail the reasons of this entanglement.
4.1.1 The more parameters, the more vulnerabilities
There is now a large literature that shows in particular that the number of parameters of a learned model increases its vulnerability, especially when it comes to privacy [KN17] and poisoning [EMGR18, Hoa24]. Intuitively, this is because the leading solution for privacy, namely differential privacy [DMNS06], requires adding a bit of noise to all parameters when updating the model, in order to correctly blur any trace of the data that led to this updating.
@@ 4.1.1 The more parameters, the more vulnerabilities
In the case of jailbreaking [GYZ+24], a key variable is the input size of the model, also known as the context window for language models. Intuitively, the larger this input size, the more the model can be tuned for specific applications via input injection (prompting in the case of language models). But then, the greater the attack surface for jailbreaking attacks [ADR+24].
4.1.2 The more generality, the more data heterogeneity, the more vulnerability
There is an additional subtler way through which the quest of generality harms resilience to attacks. Namely, to gain capabilities in a wide variety of use cases, the model needs to be trained on a highly heterogeneous set of training data. Such data are typically obtained through web crawls [Baa24], which include dubious sources [SCT23].
In particular, under such high data heterogeneity, sensitive and malicious data are significantly harder to identify and remove. More rigorously, many papers have drawn a clear connection between machine learning vulnerability and data heterogeneity [EMFG+21].
4.2 Arguments from complex system engineering
A fundamental and ubiquitous principle of complex system engineering is abstraction [Sha89]. This corresponds to hiding the complexity of the system, and to highlight instead the key features of the system as well as the limited number of ways through which it can be used. Perhaps, the most important example of such an abstraction is the creation of programming languages. Such languages allow programmers to define how they may get machines to do what they want, without having to care about the precise ways their programs will lead to transistor operations on the machines.
Abstraction is especially instrumental to implement the principles of modularisation and the “separation of concerns” [Par72, TOHSJ99]. Modularisation consists of dividing a complex systems into interacting specialised components, called modules. By carefully defining the features of the modules, which amounts to defining the abstraction they must comply with, the development of each module can then be performed independently from other modules. Similarly, each module can be evaluated, stress-tested and even correctness-proved independently from the rest of the complex system.
@@ 4.2 Arguments from complex system engineering
Modularisation is a fundamental technique of computer science that strongly highlights the value of specialisation. For instance, the hashing function SHA-256 [PVW08] is highly specialised. However, it is this amount of specialisation that has facilitated its thorough study for decades, and that has made it a core module of virtually all modern complex software systems.
@@
Better yet, clearly defining the features of each module of a complex system helps identify the module’s least required privileges [SS75]. Thereby, optimized security constraints can be enforced to the complex system’s modules. Typically, an AI model used in inference mode should not be given access to the Internet, nor to the file systems, the webcam or the keyboard. It should only be given the input data, and the required computing hardware to perform its task. Thereby, if a given module is flawed, hacked or backdoored, then the scale of the harm to the complex system will then be limited to the privileges that were given to the module. This safety oriented mindset resembles two social organisation principles: separation of powers and Subsidiarity. Separation of powers is the foundation of constitutional states and democracies [Loc90], [Mon48], and prevents the concentration of legislative, executive, and judicial powers in the hands of a single individual or political group. The separation of powers is not applied in the same way in all countries [Bel17]. The stricter it is, the more distinct, specialised, and separate the powers are, while retaining reciprocal means of action. The more flexible, the more the different powers work together. Subsidiarity is associated with decentralisation. Inspired by the social doctrine of the Catholic Church, subsidiarity [EZ14] is a political principle that assigns responsibility for a problem to the lowest competent level of public authority. This means looking for the level that is most relevant and closest to the people affected by a decision. In this way, the higher level is only called upon if the problem to be dealt with exceeds the lower level, and what can be done with the same efficiency at a lower level must be done without a higher level.
@@
Even with limited privileges, a module may still be very dangerous. This is typically the case if its output is not carefully sanitized by the subsequent algorithmic (or human) modules. For instance, if the output of a language model is used directly for email response, then the language model can be exploited to spread a spam worm [CBN24]. Likewise, a recommendation algorithm can threaten democracies by merely suggesting content, if human populations enact upon the hate speech that the algorithm amplifies. Tricking other modules can thus yield privilege escalation [ÖT24]. But crucially, by carefully specialising the modules, it will be much easier to then reason about the risks of their usage.
@@
Finally, the careful modularisation of a complex system can increase its availability by implementing the interoperable redundancy of each module, with diverse implementations of the module. The simplest example of this is the case of storage replicas. However, more inspiring systems with interoperable modules have been designed for complex tasks. Recently, the AT Protocol proposes to segment the tasks of a social media, with interoperable modules such as data storage, published message collection (know as “relays”), message labeling, feed generators and client-side applications (known as “AppViews”) [KFG+24]. By encouraging and facilitating the creation of interoperable modules managed by other entities, the protocol helps reduce the risks of social media network effects [DNW13]. Similarly, [HBC+24] proposes a modularisation of collaborative scoring, as a way to facilitate the construction of numerous interoperable subtasks.
@@
All of these elements naturally bring to mind the notion of decentralisation, studied in various fields (computer science, political sciences, management science, law, public administration, economics, etc.) with different connotations. For example, in social sciences and particularly political sciences [Tre07], it usually refers to delegation of power, transfer of skills and resources from a central power to authorities/local communities distinct from it.
@@
4.3 Arguments from economics
4.3.1 Specialisation increased competitiveness
The notion of specialisation is often associated with division of labour, which has even been a major theme in philosophy and economics for centuries ([Mar67]). In particular, division of labour should not be reduced to its technical dimension. Namely, the interdependence it generates between individuals within a society or on the scale of several has led to considering the division of labour as a important brick in the foundation of societies.
While the division and specialisation of labour within human societies can be dated back to the Neolithic period, they rather started being discussed in the 16th century. Their formalization in a systematic thinking is often associated with the writings of [Smi76], who explored the logic behind the fragmentation of work and the specialisation of tasks, as well as the links between division of labour and market competition. Smith used the example of a pin factory to illustrate the division of labour, with a work decomposed into eighteen operations required to produce one unit. He notes that each step of work fragmentation increases productivity. Unlike craftsmanship, where the craftsman controls the entire production process, manufacturing is superior in terms of productivity, due in particular to the simplification of tasks, the reduction of downtime, and control over the pace of work. He does, however, point out one limit to the division of labour: the size of the market. If the market is not big enough, there will be no outlet for the surplus production resulting from an ever-increasing division of labour.
While Smith focused on the specialisation of workers on a single production line, still within classical political economy, [Ric17] examined the division of labour on a global scale. With his theory of comparative advantage, Ricardo analysed the specialisation of national economies as they opened up to trade, using the famous example of England and Portugal and their production of wine and cloth. Ricardo’s theory has later been further developed in numerous works on the international division of labour and international specialisation ([Sam48, Jon93]).
@@ 4.3.1 Specialisation increased competitiveness
4.4 Arguments from the sociology of work, occupations, and organisations
4.4.1 Specialisation facilitates intermediate bodies
Beyond the considerations of economic competitiveness, Durkheim put forward a thesis[Dur93] asserting that the division of labour is the source of the social bond in industrial societies. He even made the study of changes in the division of labour one of the foundations of sociology. The division of labour is, he says, a source of “organic solidarity” characterized by differentiation, cooperation and strong interdependencies between individuals, as opposed to mechanical forms of social organization - individuals grouped together in communities on the basis of proximity, authority or faith. Aware of working-class misery and social conflicts, Durkheim advocated in the preface to the second edition of his book (1902) for a greater application of the division of labour to social organization, through the creation of intermediate bodies constituted on a professional basis, so that they could play the role of moral authorities arbitrating social conflicts.
@@ 4.4.1 Specialisation facilitates intermediate bodies
4.4.2 The (positive) social chain reaction of specialisation
@@
By its psychological and moral dimensions [Hug51], specialisation involves social interactions. Thus, specialised tasks arising from division of labour are part of extended and complex ensembles involving professionals, as well as non-professionals [Hug56]. Organized or not, they develop and defend (opposing) visions of what work should be. In this way, specialisation becomes part of social interactions during which legitimacy, monopolies, competition, and everything that delimits professional territories, are discussed. Specialisation is not fixed. It depends on interactions between stakeholders, and its limitations fluctuate. In the context of general AI, particularly generative AI promising to replace a considerable amount of professions and create new ones, adopting this vision of specialisation and division of labour which is not only technical, but also social, moral and psychological, would make it possible to anticipate and support emerging specialities.
4.4.3 A simplified lesson from company towns
@@
In an era where AGI is increasingly becoming the proclaimed goal of AI development, and when the most powerful billionaire are talking about the “everything app”[Bel24], it might be worth considering the 19th and 20th century use cases of company towns [Gar92]. As defined in Wikipedia, “a company town is a place where all or most of the stores and housing in the town are owned by the same company that is also the main employer. Company towns are often planned with a suite of amenities such as stores, houses of worship, schools, markets, and recreation facilities”. As a non-specialised, generalist, non-human entity; or as a general intelligence to stay in the parlance of AI, company towns can be seen as dealing with needs which are not limited to the sphere of work (in this case, postal, school, health, food, etc. services). Linked to what is known as “industrial paternalism” [Noi88], this particular type of work management is associated with organisational problems. For example, the inability to provide all the services promised to workers, given that in fact, this system designed for blue-collar workers, mainly offered its advantages to white-collar workers. Another example is the difficulty of supervising workers, because it is impossible to be efficient by playing the roles of employer, priest, carer, policeman, etc., all at the same time, and therefore to bypass the division of labour and specialisation. Over and above the untenable nature of this strict social control of workers for reasons of labour profitability, this model of work organisation, based on task and power concentration, has not lasted. The reasons for the collapse of this system arguably include the replacement of the State by the management of the company town, the replacement of national law by the company practices, and most importantly, the replacement of the centuries-long process of specialisation of tasks and activities, by the one-company-do-it-all mindset of company-towns.
@@ 5.1 Documentations
In the context of machine learning, increased calls for documentation have arisen, especially in the case of machine learning models through model cards [MWZ+19] and in the case of datasets through data sheets [GMV+21].
5.2 Type systems and proofs of correctness
However, careful specification can yield much more significant improvements, especially with regards to the resilience and security of complex systems. In particular, modern programming languages provide sophisticated type systems [Car96, MK14, GSJ+24], which allow to directly encode the specifications of a module, and mathematically prove that the module correctly implements the specifications at compilation time. Crucially, this allows to catch bugs before deployment.
@@ 5.2 Type systems and proofs of correctness
Now, it is noteworthy that type systems allow verification at compilation time, given the source code. In practice, an additional challenge is to verify the correctness of a program, given its compiled binary code. To meet this challenge, the fascinating domain of verifiable computing [AWH+18]. In addition to multi-party computations [GS22], where carefully specified operations and communications enable more secure information processing, verifiable computing leverages powerful cryptographic primitives like Succinct Non-Interactive Arguments of Knowledge (SNARK) [T+22], which enable a powerful computer to prove the soundness of its output to a weak verifier.
Recently, there has in fact been much progress towards verifiable machine-learning computing, at inference time [FMZ+24a] (after deployment) and at training time [FMZ+24b] (before deployment). In particular, these solutions allow to envisage the effortless enforcement of some existing laws (like EU’s AI Act), e.g. by demanding that all commercialized AI systems prove that they were trained on legally obtained data (i.e. excluding copyrighted, sensitive or error-prone data). In fact, remarkably, the proofs may be constructed in zero-knowledge, i.e. without disclosing non-legally-binding proprietary data.
@@
5.3 Hypertelia and the pitfalls of automated task specification
In biology, hypertelia [BVW74] designates an exaggerated growth of certain organs in relation to their function, to the point of making them annoying for the animal and its entire species (cf. canines of the saber-toothed tiger, too heavy antlers of the deer or disproportionate tail of the peacock). Taken up in philosophy by Simondon [Sim58, SMH80] for his analysis of technical objects, the notion expresses in this context the idea of an exaggerated specialisation, and a functional over-adaptation of technical objects. Adapting an object too much to a particular purpose and context can result in its inability to function well. Tools adjusted to very specialized circumstances can lose autonomy outside of their specific technical environment. Specialisation must be guided by proven mechanisms and not be an end in itself. If guided by human choices, these can be inspired by the specialisation that has taken place in the professions. The challenge of good specification will arise more intensely when the granularity of specialisations is finer, which is enabled by automation. This is where a risk of hypertelia can occur, without being easily relieved. The challenge here is to determine who or what meta-algorithm defines the speciality of each sub-algorithm.
5.4 Two limits of specifications
@@
First, some tasks are too complex to specify, in the sense of the Kolomogorov-Solomonoff complexity [Sol60, Kol63]. Formally, a formalized specification is a program that, given any program, returns whether the program verifies the specification. The Kolomogorov-Solomonoff complexity is defined as the shortest program that implements this formalized specification. Unfortunately, it is conjectured that many tasks require extremely complex specifications, in the sense that they cannot be formally described in less than a million lines of codes. As an example, EU’s AI Act is 144 pages long. Yet, it is clearly far from formalized and has been argued to be very incomplete [LWM24]. Another example is today’s web standards, which correspond to very long documents that specify languages like HTML, CSS and ECMAScript, among others.
Second, some specifications are extremely hard to agree on. This is typically the case of content moderation, but it also holds for text autocompletion, biasless image generation or content recommendation. Given additionally the lack of knowledge about the human population’s distribution of specification preferences, and perhaps even about one’s own preferred specification, a choice of specification may seem premature and inappropriate. This is one heart of the challenges surrounding “AI alignment” [HEM19].
@@
5.5 Specifying unspecifiable tasks through specified governance
@@
Remarkably, a growing line of research has provided new solutions for algorithmic governance. In particular, WeBuildAI [LKK+19] proposed a software which allow a number of stakeholders to collaboratively select the recipients of food donations. This software was itself subject to both informal and formal specifications, e.g. both donors, recipients, volunteers and the organising nonprofit association should all have some voting power on the decision, or they may either use the software’s learning-based system to construct a model of their preferences, or write themselves a model of their preferences. This system has subsequently inspired other algorithmic governance systems, e.g. for trolley dilemmas [NGA+18], kidney donation [FBSA+20] content recommendation [HFJ+21], proposal prioritization [SBE+21] and contextual note selection [RSDP23].
In fact, [HBC+24] aims to clarify the challenges of the specification of the algorithmic collaborative governance of the scoring of any set of alternatives. The list of subtasks to better specify includes participant verification, trust propagation through a web of trust, preference generalization, and secure preference aggregation, among others.

### v2

@@ 4 Why Specialise?
Large Models are More Vulnerable
In the context of artificial intelligence, generality is obtained by training large models on massive web data crawls. Indeed, ever larger models and datasets have been empirically observed to yield ever more spectacular performances (Brown et al. 2020; Kaplan et al. 2020b; Dubey et al. 2024). However, these high performances seem inevitably entangled with a number of security issues (El-Mhamdi et al. 2022; Oprea and Vassilev 2023), like privacy violations, jailbreaking and data poisoning. Below we detail the reasons of this entanglement.
The more parameters, the more vulnerabilities
There is now a large literature that shows in particular that the number of parameters of a learned model increases its vulnerability, especially when it comes to privacy (Kattis and Nikolov 2017) and poisoning (El-Mhamdi et al. 2018; Hoang 2024; El-Mhamdi et al. 2017). Intuitively, this is because the leading solution for privacy, namely differential privacy (Dwork et al. 2006), requires adding a bit of noise to all parameters when updating the model, in order to correctly blur any trace of the data that led to this updating.
@@
In the case of jailbreaking (Guo et al. 2024), a key variable is the input size of the model, also known as the context window for language models. Intuitively, the larger this input size, the more the model can be tuned for specific applications via input injection (prompting in the case of language models). But then, the greater the attack surface for jailbreaking attacks (Anil et al. 2024).
The more generality, the more data heterogeneity, the more vulnerability
There is an additional subtler way through which the quest of generality harms resilience to attacks. Namely, to gain capabilities in a wide variety of use cases, the model needs to be trained on a highly heterogeneous set of training data. Such data are typically obtained through web crawls (Baack 2024), which include dubious sources (Schaul, Chen, and Tiku 2023).
In particular, under such high data heterogeneity, sensitive and malicious data are significantly harder to identify and remove. More rigorously, many papers have drawn a clear connection between machine learning vulnerability and data heterogeneity (El-Mhamdi et al. 2021).
Arguments from Complex System Engineering
A fundamental and ubiquitous principle of complex system engineering is abstraction (Shaw 1989). This corresponds to hiding the complexity of the system, and to highlight instead the key features of the system as well as the limited number of ways through which it can be used. Perhaps, the most important example of such an abstraction is the creation of programming languages. Such languages allow programmers to define how they may get machines to do what they want, without having to care about the precise ways their programs will lead to transistor operations on the machines.
Abstraction is especially instrumental to implement the principles of modularisation and the “separation of concerns” (Parnas 1972; Tarr et al. 1999). Modularisation consists of dividing a complex systems into interacting specialised components, called modules. By carefully defining the features of the modules, which amounts to defining the abstraction they must comply with, the development of each module can then be performed independently from other modules. Similarly, each module can be evaluated, stress-tested and even correctness-proved independently from the rest of the complex system.
@@
Modularisation is a fundamental technique of computer science that strongly highlights the value of specialisation. For instance, the hashing function SHA-256 (Penard and Van Werkhoven 2008) is highly specialised. However, it is this amount of specialisation that has facilitated its thorough study for decades, and that has made it a core module of virtually all modern complex software systems.
@@
Better yet, clearly defining the features of each module of a complex system helps identify the module’s least required privileges (Saltzer and Schroeder 1975). Thereby, optimized security constraints can be enforced to the complex system’s modules. Typically, an AI model used in inference mode should not be given access to the Internet, nor to the file systems, the webcam or the keyboard. It should only be given the input data, and the required computing hardware to perform its task. Thereby, if a given module is flawed, hacked or backdoored, then the scale of the harm to the complex system will then be limited to the privileges that were given to the module. This safety oriented mindset resembles two social organisation principles: separation of powers and Subsidiarity. Separation of powers is the foundation of constitutional states and democracies (Locke 1690), (Montesquieu 1748), and prevents the concentration of legislative, executive, and judicial powers in the hands of a single individual or political group. The separation of powers is not applied in the same way in all countries (Bellamy 2017). The stricter it is, the more distinct, specialised, and separate the powers are, while retaining reciprocal means of action. The more flexible, the more the different powers work together. Subsidiarity is associated with decentralisation. Inspired by the social doctrine of the Catholic Church, subsidiarity (Evans and Zimmermann 2014) is a political principle that assigns responsibility for a problem to the lowest competent level of public authority. This means looking for the level that is most relevant and closest to the people affected by a decision. In this way, the higher level is only called upon if the problem to be dealt with exceeds the lower level, and what can be done with the same efficiency at a lower level must be done without a higher level.
@@
Even with limited privileges, a module may still be very dangerous. This is typically the case if its output is not carefully sanitized by the subsequent algorithmic (or human) modules. For instance, if the output of a language model is used directly for email response, then the language model can be exploited to spread a spam worm (Cohen, Bitton, and Nassi 2024). Likewise, a recommendation algorithm can threaten democracies by merely suggesting content, if human populations enact upon the hate speech that the algorithm amplifies. Tricking other modules can thus yield privilege escalation (Özdemir and Tuna 2024). But crucially, by carefully specialising the modules, it will be much easier to then reason about the risks of their usage.
@@
Finally, the careful modularisation of a complex system can increase its availability by implementing the interoperable redundancy of each module, with diverse implementations of the module. The simplest example of this is the case of storage replicas. However, more inspiring systems with interoperable modules have been designed for complex tasks. Recently, the AT Protocol proposes to segment the tasks of a social media, with interoperable modules such as data storage, published message collection (know as “relays”), message labeling, feed generators and client-side applications (known as “AppViews”) (Kleppmann et al. 2024). By encouraging and facilitating the creation of interoperable modules managed by other entities, the protocol helps reduce the risks of social media network effects (Dou, Niculescu, and Wu 2013). Similarly, (Hoang et al. 2024) proposes a modularisation of collaborative scoring, as a way to facilitate the construction of numerous interoperable subtasks.
@@
All of these elements naturally bring to mind the notion of decentralisation, studied in various fields (computer science, political sciences, management science, law, public administration, economics, etc.) with different connotations. For example, in social sciences and particularly political sciences (Treisman 2007), it usually refers to delegation of power, transfer of skills and resources from a central power to authorities/local communities distinct from it.
@@
Arguments from Economics
Specialisation Increases Competitiveness.
The notion of specialisation is often associated with division of labour, which has even been a major theme in philosophy and economics for centuries ((Marx 1867)). In particular, division of labour should not be reduced to its technical dimension. Namely, the interdependence it generates between individuals within a society or on the scale of several has led to considering the division of labour as a important brick in the foundation of societies.
While the division and specialisation of labour within human societies can be dated back to the Neolithic period, they rather started being discussed in the 16th century. Their formalization in a systematic thinking is often associated with the writings of (Smith 1776), who explored the logic behind the fragmentation of work and the specialisation of tasks, as well as the links between division of labour and market competition. Smith used the example of a pin factory to illustrate the division of labour, with a work decomposed into eighteen operations required to produce one unit. He notes that each step of work fragmentation increases productivity. Unlike craftsmanship, where the craftsman controls the entire production process, manufacturing is superior in terms of productivity, due in particular to the simplification of tasks, the reduction of downtime, and control over the pace of work. He does, however, point out one limit to the division of labour: the size of the market. If the market is not big enough, there will be no outlet for the surplus production resulting from an ever-increasing division of labour.
While Smith focused on the specialisation of workers on a single production line, still within classical political economy, (Ricardo 1817) examined the division of labour on a global scale. With his theory of comparative advantage, Ricardo analysed the specialisation of national economies as they opened up to trade, using the famous example of England and Portugal and their production of wine and cloth. Ricardo’s theory has later been further developed in numerous works on the international division of labour and international specialisation ((Samuelson 1948; Jones 1993)).
@@
Arguments from the Sociology of Work, Occupations, and Organisations
Specialisation enables intermediate bodies.
Beyond the considerations of economic competitiveness, Durkheim put forward a thesis(Durkheim 1893) asserting that the division of labour is the source of the social bond in industrial societies. He even made the study of changes in the division of labour one of the foundations of sociology. The division of labour is, he says, a source of “organic solidarity” characterized by differentiation, cooperation and strong interdependencies between individuals, as opposed to mechanical forms of social organization - individuals grouped together in communities on the basis of proximity, authority or faith. Aware of working-class misery and social conflicts, Durkheim advocated in the preface to the second edition of his book (1902) for a greater application of the division of labour to social organization, through the creation of intermediate bodies constituted on a professional basis, so that they could play the role of moral authorities arbitrating social conflicts.
@@
The (positive) social chain reaction of specialisation.
@@
By its psychological and moral dimensions (Hughes 1951), specialisation involves social interactions. Thus, specialised tasks arising from division of labour are part of extended and complex ensembles involving professionals, as well as non-professionals (Hughes 1956). Organized or not, they develop and defend (opposing) visions of what work should be. In this way, specialisation becomes part of social interactions during which legitimacy, monopolies, competition, and everything that delimits professional territories, are discussed. Specialisation is not fixed. It depends on interactions between stakeholders, and its limitations fluctuate. In the context of general AI, particularly generative AI promising to replace a considerable amount of professions and create new ones, adopting this vision of specialisation and division of labour which is not only technical, but also social, moral and psychological, would make it possible to anticipate and support emerging specialities.
A simplified lesson from company towns.
@@
In an era where AGI is increasingly becoming the proclaimed goal of AI development, and when the most powerful billionaire are talking about the “everything app”(Belanger 2024), it might be worth considering the 19th and 20th century use cases of company towns (Garner 1992). As defined in Wikipedia, “a company town is a place where all or most of the stores and housing in the town are owned by the same company that is also the main employer. Company towns are often planned with a suite of amenities such as stores, houses of worship, schools, markets, and recreation facilities”. As a non-specialised, generalist, non-human entity; or as a general intelligence to stay in the parlance of AI, company towns can be seen as dealing with needs which are not limited to the sphere of work (in this case, postal, school, health, food, etc. services). Linked to what is known as “industrial paternalism” (Noiriel 1988), this particular type of work management is associated with organisational problems. For example, the inability to provide all the services promised to workers, given that in fact, this system designed for blue-collar workers, mainly offered its advantages to white-collar workers. Another example is the difficulty of supervising workers, because it is impossible to be efficient by playing the roles of employer, priest, carer, policeman, etc., all at the same time, and therefore to bypass the division of labour and specialisation. Over and above the untenable nature of this strict social control of workers for reasons of labour profitability, this model of work organisation, based on task and power concentration, has not lasted. The reasons for the collapse of this system arguably include the replacement of the State by the management of the company town, the replacement of national law by the company practices, and most importantly, the replacement of the centuries-long process of specialisation of tasks and activities, by the one-company-do-it-all mindset of company-towns.
@@ 5 A Case for Specification
In the context of machine learning, increased calls for documentation have arisen, especially in the case of machine learning models through model cards (Mitchell et al. 2019) and in the case of datasets through data sheets (Gebru et al. 2021).
Type Systems and Proofs of Correctness
However, careful specification can yield much more significant improvements, especially with regards to the resilience and security of complex systems. In particular, modern programming languages provide sophisticated type systems (Cardelli 1996; Matsakis and Klock 2014; Gäher et al. 2024), which allow to directly encode the specifications of a module, and mathematically prove that the module correctly implements the specifications at compilation time. Crucially, this allows to catch bugs before deployment.
@@
Now, it is noteworthy that type systems allow verification at compilation time, given the source code. In practice, an additional challenge is to verify the correctness of a program, given its compiled binary code. To meet this challenge, the fascinating domain of verifiable computing (Ahmad et al. 2018). In addition to multi-party computations (Garg and Srinivasan 2022), where carefully specified operations and communications enable more secure information processing, verifiable computing leverages powerful cryptographic primitives like Succinct Non-Interactive Arguments of Knowledge (SNARK) (Thaler et al. 2022), which enable a powerful computer to prove the soundness of its output to a weak verifier.
Recently, there has in fact been much progress towards verifiable machine-learning computing, at inference time (Fan et al. 2024a) (after deployment) and at training time (Fan et al. 2024b) (before deployment). In particular, these solutions allow to envisage the effortless enforcement of some existing laws (like EU’s AI Act), e.g. by demanding that all commercialized AI systems prove that they were trained on legally obtained data (i.e. excluding copyrighted, sensitive or error-prone data). In fact, remarkably, the proofs may be constructed in zero-knowledge, i.e. without disclosing non-legally-binding proprietary data.
@@
Hypertelia and the Pitfalls of Automated Task Specification
In biology, hypertelia (Brunner Von Wattenwyl 1874) designates an exaggerated growth of certain organs in relation to their function, to the point of making them annoying for the animal and its entire species (cf. canines of the saber-toothed tiger, too heavy antlers of the deer or disproportionate tail of the peacock). Taken up in philosophy by Simondon (Simondon 1958; Simondon, Mellamphy, and Hart 1980) for his analysis of technical objects, the notion expresses in this context the idea of an exaggerated specialisation, and a functional over-adaptation of technical objects. Adapting an object too much to a particular purpose and context can result in its inability to function well. Tools adjusted to very specialized circumstances can lose autonomy outside of their specific technical environment. Specialisation must be guided by proven mechanisms and not be an end in itself. If guided by human choices, these can be inspired by the specialisation that has taken place in the professions. The challenge of good specification will arise more intensely when the granularity of specialisations is finer, which is enabled by automation. This is where a risk of hypertelia can occur, without being easily relieved. The challenge here is to determine who or what meta-algorithm defines the speciality of each sub-algorithm.
Two Limits of Specifications
@@
First, some tasks are too complex to specify, in the sense of the Kolomogorov-Solomonoff complexity (Solomonoff 1960; Kolmogorov 1963). Formally, a formalized specification is a program that, given any program, returns whether the program verifies the specification. The Kolomogorov-Solomonoff complexity is defined as the shortest program that implements this formalized specification. Unfortunately, it is conjectured that many tasks require extremely complex specifications, in the sense that they cannot be formally described in less than a million lines of codes. As an example, EU’s AI Act is 144 pages long. Yet, it is clearly far from formalized and has been argued to be very incomplete (Laux, Wachter, and Mittelstadt 2024). Another example is today’s web standards, which correspond to very long documents that specify languages like HTML, CSS and ECMAScript, among others.
Second, some specifications are extremely hard to agree on. This is typically the case of content moderation, but it also holds for text autocompletion, biasless image generation or content recommendation. Given additionally the lack of knowledge about the human population’s distribution of specification preferences, and perhaps even about one’s own preferred specification, a choice of specification may seem premature and inappropriate. This is one heart of the challenges surrounding “AI alignment” (Hoang and El Mhamdi 2019; El-Mhamdi and Hoang 2024; Majka and El-Mhamdi 2025).
@@
Specifying Unspecifiable Tasks Through Specified Governance
@@
Remarkably, a growing line of research has provided new solutions for algorithmic governance. In particular, WeBuildAI (Lee et al. 2019) proposed a software which allow a number of stakeholders to collaboratively select the recipients of food donations. This software was itself subject to both informal and formal specifications, e.g. both donors, recipients, volunteers and the organising nonprofit association should all have some voting power on the decision, or they may either use the software’s learning-based system to construct a model of their preferences, or write themselves a model of their preferences. This system has subsequently inspired other algorithmic governance systems, e.g. for trolley dilemmas (Noothigattu et al. 2018), kidney donation (Freedman et al. 2020) content recommendation (Hoang et al. 2021), proposal prioritization (Small et al. 2021) and contextual note selection (Righes et al. 2023). More recently, (Hoang et al. 2024) aims to clarify the challenges of the specification of the algorithmic collaborative governance of the scoring of any set of alternatives. The list of subtasks to better specify includes participant verification, trust propagation through a web of trust, preference generalization, and secure preference aggregation, among others.